OpenStudy (anonymous):

Simplify (3√2 +√3)(√3 -√2)

OpenStudy (mathstudent55):

Use FOIL to multiply it out

OpenStudy (anonymous):

can you help me

OpenStudy (anonymous):

ya nitz your here

OpenStudy (anonymous):

\[3\sqrt{2}(\sqrt{3}-\sqrt{2})+\sqrt{3}(\sqrt{3}-\sqrt{2})\]

OpenStudy (anonymous):

\[3\sqrt 6-6+3-\sqrt{6}\]

OpenStudy (anonymous):

\[2\sqrt{6}-3\]

OpenStudy (anonymous):

Thabks

OpenStudy (anonymous):

*thanks

OpenStudy (anonymous):

ANY OTHER QUERY